### v1

@@ 3.5 Data driven formulation
The surrogate flux map $\tilde{Z}:L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))\rightarrow L^{2}(S,\nu_{M};\mathbb{R}^{d})$ maps the operator field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L})))$ to the flux field $\tilde{z}\in L^{2}(S,\nu_{M};\mathbb{R}^{d})$, $d\in\mathbb{N}$, where each vector of the latter corresponds to the computed surrogate flux values for each sampled gridpoint. More specifically, given $d$ quantum observables $A^{(j)}\in B(H_{\Omega,N}^{L})$, $j\in\{0,\ldots,d-1\}$, and a quantum state field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$, each component of the flux map $\tilde{Z}$ acts by $\tilde{Z}^{(j)}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A^{(j)})\in\mathbb{R}$. By slightly abusing notation we may write more compactly $\tilde{z}_{m}^{(j)}=\tilde{Z}^{(j)}(\rho_{m})=\operatorname{tr}(\rho_{m}A^{(j)})$. Using our basis representation of each $A^{(j)}$ by matrix $\bm{A}^{(j)}\in\mathbb{R}^{L\times L}$ and $\rho_{m}$ by vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$, we have that $\tilde{z}_{m}^{(j)}=\underline{\rho}_{m}^{T}\bm{A}^{(j)}\underline{\rho}_{m}\in\mathbb{R}$ where superscript $T$ denotes transposition.
@@
The induced transfer operator is defined as $\mathcal{P}_{N}^{L}:B(H_{\Omega,N}^{L})\rightarrow B(H_{\Omega,N}^{L})$. Because we work with pure quantum states, which can be represented by vectors instead of matrices, it is sufficient to approximate transfer operator $P_{N}^{L}:H_{\Omega,N}^{L}\rightarrow H_{\Omega,N}^{L}$ instead of $\mathcal{P}_{N}^{L}$. To build $P_{N}^{L}$ we begin by approximating $P_{N}:H_{\Omega,N}\rightarrow H_{\Omega,N}$ by the shift operator $\hat{P}_{N}:H_{\Omega,N}\rightarrow H_{\Omega,N}$
$\hat{P}_{N}f(\hat{x}_{n},s_{m})=\begin{cases}f(\hat{x}_{n+1},s_{m})&\text{if $0\leq n<N-1$}\\ 0&\text{if $n=N-1$}.\end{cases}$
@@
After projection we obtain $P_{N}^{L}=\operatorname{Proj}_{L}\hat{P}_{N}$ represented by a matrix $\bm{P}_{N}^{L}\in\mathbb{R}^{L\times L}$, $(P_{N}^{L})_{ij}=\langle\phi_{i},P_{N}^{L}\phi_{j}\rangle_{N}$. Given a single quantum state $\rho_{i}\in Q(H_{\Omega,N}^{L})$ at timestep $t_{i}$ with vector representation $\underline{\rho}_{i}\in\mathbb{R}^{L}$, we obtain the updated state $\underline{\rho}_{i+1}$ by
$\underline{\rho}_{i+1}=\frac{(\bm{P}_{N}^{L})^{T}\underline{\rho}_{i}}{\lVert(\bm{P}_{N}^{L})^{T}\underline{\rho}_{i}\rVert_{2}}$
@@
where superscript $T$ denotes transposition. Matrix $\bm{P}_{N}^{L}$ is in general not unitary, which is why we renormalize the updated quantum state after applying $\bm{P}_{N}^{L}$.
@@
Quantum effects are operators $e\in B(H_{\Omega,N}^{L})$ used to condition quantum density states based on a given state vector $\omega\in\Omega$. To build effect $e$ from state vector $\omega$ we introduce kernel function $k_{c}:\Omega\times\Omega\rightarrow\mathbb{R}$, defined as the pullback of kernel $k_{cw}:\mathcal{W}_{c}\times\mathcal{W}_{c}\rightarrow\mathbb{R}$, $k_{c}(\omega,\omega^{\prime})=k_{cw}(W_{c}(\omega),W_{c}(\omega^{\prime}))$ with data space $\mathcal{W}_{c}$ and associated data map $W_{c}:\Omega\rightarrow\mathcal{W}_{c}$. To define kernel function $k_{cw}$ we define the data space $\mathcal{W}_{c}=\mathbb{R}^{J}$ with parameter $J\in\mathbb{N}$. For a given gridpoint $s_{m}\in S_{M}$ and state vector $\omega_{m}=(x,s_{m})\in\Omega$ we define the action of data map $W_{J}:\Omega\rightarrow\mathbb{R}^{J}$ as
$W_{J}((x,s_{m}))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(s_{m-(J-1)/2}),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(s_{m}),\ldots,\\ \operatorname{proj}_{\mathcal{X}}(x)(s_{m+(J-1)/2})\bigr{)}.$
The kernel $k_{cw}:\mathbb{R}^{J}\times\mathbb{R}^{J}\rightarrow\mathbb{R}$ is then defined as
@@
(9) $k_{cw}(W_{J}(\omega),W_{J}(\omega^{\prime}))=\exp\Bigl{(}-\frac{1}{\epsilon^{\prime}J}\lVert W_{J}(\omega)-W_{J}(\omega^{\prime})\rVert_{J}^{2}\Bigr{)}$
@@
For a fixed gridpoint $s_{m}\in S_{M}$, given a state vector $\omega_{m}$ and the full training dataset $\{\omega_{nm}\}_{nm=0}^{NM-1}$, kernel $k_{c}$ is used to produce function $f_{m}\in H_{\Omega,N}$, $f_{m}(\cdot)=k_{c}(\omega_{m},\cdot)$, leading to multiplication operator $M_{f_{m}}$ and the desired quantum effect $e_{m}=\operatorname{Proj}_{L}M_{f_{m}}$. Employing the quantum Bayes rule (2) then yields the conditioned quantum state $\rho_{m}|_{e_{m}}$ for gridpoint $s_{m}$. Intuitively, function $f_{m}$ can be thought of as an affinity function on state space $\Omega$, measuring how “similar” the given state $\omega_{m}$ is to the states included in our training dataset. Building operator $e_{m}$ and conditioning the prior state $\rho_{m}$ then yields the posterior state $\rho_{m}|_{e_{m}}$ that encodes the state space affinities captured by $f_{m}$. We often refer to $f_{m}$ as the feature vector produced by kernel $k_{c}$ for product state vector $\omega_{m}$.
With respect to our basis representation of $\rho_{m}$ by vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$ and effect $e_{m}$ by matrix $\bm{e}_{m}\in\mathbb{R}^{L\times L}$, the conditioning operation takes the form
@@
To avoid having to compute the square root of matrix $\bm{e}_{m}$, we modify the way we form effect $e_{m}$. We use function $f_{m}^{1/2}\in H_{\Omega,N}$ to form multiplication operator $\tilde{A}_{m}=M_{f_{m}}^{1/2}$ and define effect $\tilde{e}_{m}=(\operatorname{Proj}_{L}\tilde{A}_{m})^{2}=(\tilde{A}_{m}^{L})^{2}$. Conditioning with respect to $\tilde{e}_{m}$ now takes the form
@@
Although effects $\tilde{e}_{m}$ and $e_{m}$ are generally different, the two are equal when $L=N$, and thus also share the same asymptotic limit as $N\rightarrow\infty$ and $L\rightarrow\infty$ [17].
The data map $W_{J}$ embeds state vector $(x,s_{m})\in\Omega$ into $\mathbb{R}^{J}$ by considering the resolved state values at gridpoints neighboring $s_{m}\in S_{M}$, mimicking the structure of the stencil used for the spatial discretization of the considered dynamics. In this way kernel $k_{cw}$ compares the stencil of values around $s_{m}$ to the corresponding stencil of points for every state included in our training data. This makes for a stronger affinity condition than comparing values of states at only a single gridpoint. For hyperbolic partial differential equations, where the notion of a local propagation velocity is well defined, using a stencil of values is analogous to using a time delayed signal with a generally fractional number of delays.
Because the spatially local value $\omega_{m}=(x,s_{m})$ is compared to every other state vector present in our training data $\{\omega_{nm}\}_{nm=0}^{NM-1}$, spatial translations of state $x$ are naturally factored during the conditioning process. This means that if a spatial translation of state $x$ is present in our training dataset, then the similarity between product state $\omega_{m}=(x,s_{m})$ and the corresponding value of the spatial translation of state $x$ will be identified in the conditioning process. For partial differential equations with spatial translation symmetry, this means that we do not have to artificially inflate our training dataset by including all spatial translations of the sampled states.
@@ 3.6 Multitrajectory data
To build the data driven formulation of the scheme we have so far relied on a single dynamical trajectory $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$. In this section we generalize this construction to the case where our training data consists of multiple trajectories.
We consider $K\in\mathbb{N}$ trajectories of $N_{k}$ samples each, denoted by $\{x_{n}^{(k)}\}_{n=0}^{N_{k}-1}$, $k\in\{0,\ldots,K-1\}$. Considering each trajectory independently we use the approach outlined in Section 3.5 to build an orthonormal basis for the Hilbert space $H_{\Omega,N_{k}}$, which is the space of functions represented by their values on the samples of the $k$th trajectory only. We use the $L_{k}$ leading eigenfunctions $\{\phi_{\ell}^{(k)}\}_{\ell=0}^{L_{k}-1}$ to define subspace $H_{\Omega,N_{k}}^{L_{k}}\subset H_{\Omega,N_{k}}$. Denoting $N=\sum_{k=0}^{K-1}N_{k}$ and $L=\sum_{k=0}^{K-1}L_{k}$ we define the combined space
$H_{\Omega,N}^{L}=H_{\Omega,N_{0}}^{L_{0}}\oplus\cdots\oplus H_{\Omega,N_{K-1}}^{L_{K-1}},$
with the direct sum of the individual bases producing a basis for the combined space $H_{\Omega,N}^{L}$. The basis derived in this way consists of a total of $L$ eigenfunctions, each representing the values of the function on $N$ samples. For each eigenfunction, all of its values are zero except for the block of values that corresponds to the $N_{k}$ samples of the $k$th trajectory it was derived from.
Having derived the basis for space $H_{\Omega,N}^{L}$, the construction of the rest of the scheme proceeds as in Section 3.5. The only additional difference is in the approximation of the transfer operator $P_{N}$ by the shift operator $\hat{P}_{N}$, where time shifts are now applied only within each trajectory, and the last temporal sample of each trajectory is mapped to zero. Namely, we make sure not to shift samples across different trajectories.
@@
Using a single dynamical trajectory is generally sufficient when considering a dynamical system possessing a unique physical measure, namely a measure that can be sampled by a set of initial conditions of positive measure with respect to the ambient measure. These are measures whose sampling can be realized numerically, with SRB measures being an important example [43]. In such systems, a single long trajectory can be used to sample the physical measure of the dynamics.
@@
Using multiple trajectories as training data is of interest when considering dynamical systems possessing multiple physical measures. For instance, hamiltonian systems or hyperbolic systems with several competing basins of attraction. Additionally, using multiple trajectories is necessary in parametric forecasting, where one considers a dynamical system whose qualitative dynamics changes with respect to a bifurcation parameter.
@@
Finally, even in systems with a unique physical measure, being able to use multiple trajectories to sample that measure allows one to consider optimal sampling strategies. In designing such a strategy, one tries to optimize the sampling rate with respect to a chosen objective, instead of relying on the inherent rate provided by a single dynamical trajectory.

### v2

@@ 4.4 Data driven formulation
For each $j\in\{0,\ldots,d-1\}$ the surrogate flux function $\tilde{z}_{j}\colon L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))\to H_{S}$ maps the quantum state field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L})))$ to the scalar field $\tilde{z}_{j}(\rho)\in H_{S}$ with $\tilde{z}_{j}(\rho)(s_{m})\in\mathbb{R}$ the computed flux at gridpoint $s_{m}\in S_{M}$. More specifically, given $d$ quantum observables $A_{j}\in B(H_{\Omega,N}^{L})$ and a quantum state field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ each surrogate flux function $\tilde{z}_{j}$ acts by $\tilde{z}_{j}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A_{j})\in\mathbb{R}$. Using our basis representation of each $A_{j}$ by matrix $\bm{A}_{j}\in\mathbb{R}^{L\times L}$ and $\rho(s_{m})$ by vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$ we have that $\tilde{z}_{j}(\rho)(s_{m})=\underline{\rho}_{m}^{\top}\bm{A}_{j}\underline{\rho}_{m}$.
@@
The evolution of quantum states is governed by the projected induced transfer operator $\mathcal{P}_{N}^{L}\colon B(H_{\Omega,N}^{L})\to B(H_{\Omega,N}^{L})$. Because we work with pure quantum states—represented by vectors instead of matrices—we are going to form the projected transfer operator $P_{N}^{L}\colon H_{\Omega,N}^{L}\to H_{\Omega,N}^{L}$ instead of $\mathcal{P}_{N}^{L}$. To build $P_{N}^{L}$ we define the shift operator $P_{N}\colon H_{\Omega,N}\to H_{\Omega,N}$
$P_{N}f(\hat{x}_{n},s_{m})=\begin{cases}0&\text{if $n=0$}\\ f(\hat{x}_{n-1},s_{m})&\text{if $1\leq n<N$}.\end{cases}$
@@
After projection we obtain $P_{N}^{L}=\operatorname{Proj}_{L}P_{N}$ represented by a matrix $\bm{P}_{N}^{L}\in\mathbb{R}^{L\times L}$, $(P_{N}^{L})_{ik}=\langle\phi_{i},P_{N}^{L}\phi_{k}\rangle_{N}$. Given a single quantum state $\rho_{n}\in Q(H_{\Omega,N}^{L})$ at time $t_{n}$ with vector representation $\underline{\rho}_{n}\in\mathbb{R}^{L}$ we obtain the updated state $\underline{\rho}_{n+1}$ by
$\underline{\rho}_{n+1}=\frac{\bm{P}_{N}^{L}\,\underline{\rho}_{n}}{\lVert\bm{P}_{N}^{L}\,\underline{\rho}_{n}\rVert_{2}}.$
@@
The matrix $\bm{P}_{N}^{L}$ is in general not unitary, which is why we renormalize the updated quantum state after applying $\bm{P}_{N}^{L}$.
@@
Next, quantum effects are operators $e\in B(H_{\Omega,N}^{L})$ used to condition quantum states based on a given state vector $\omega\in\Omega$. To build effect $e$ from state vector $\omega$ we introduce the conditioning kernel function $\kappa_{c}\colon\Omega\times\Omega\to\mathbb{R}$, defined as the pullback of kernel $k_{c}\colon\mathcal{W}_{c}\times\mathcal{W}_{c}\to\mathbb{R}$, $\kappa_{c}(\omega,\omega^{\prime})=k_{c}(W_{c}(\omega),W_{c}(\omega^{\prime}))$ with conditioning data space $\mathcal{W}_{c}$ and associated data map $W_{c}\colon\Omega\to\mathcal{W}_{c}$. To define kernel $k_{c}$ we introduce the data space $\mathcal{W}_{c}=\mathbb{R}^{J}$ with parameter $J\in\mathbb{N}$. Additionally, we define the neighborhood function $\mathcal{N}\colon S\to S^{J}$ which for a given gridpoint $s_{m}\in S_{M}$ returns its $J$ nearest neighbors on $S_{M}$ in a specified order. For a given gridpoint $s_{m}\in S_{M}$ and state vector $\omega_{m}=(x,s_{m})\in\Omega$ we then define the action of data map $W_{J}\colon\Omega\to\mathbb{R}^{J}$ by
$W_{J}((x,s_{m}))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(\mathcal{N}^{(0)}(s_{m})),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(\mathcal{N}^{(J-1)}(s_{m}))\bigr{)}$
where $\mathcal{N}^{(j)}(s_{m})$ denotes the $j$th component of vector $\mathcal{N}(s_{m})$. For example, for a one dimensional grid $S_{M}$ and odd $J\in\mathbb{N}$ the action of $W_{J}$ becomes
$W_{J}((x,s_{m}))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(s_{m-(J-1)/2}),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(s_{m}),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(s_{m+(J-1)/2})\bigr{)}.$
The kernel $k_{c}\colon\mathbb{R}^{J}\times\mathbb{R}^{J}\to\mathbb{R}$ is then defined as
@@
$k_{c}(W_{J}(\omega),W_{J}(\omega^{\prime}))=\exp\bigl{(}-\frac{1}{\epsilon^{\prime}J}\lVert W_{J}(\omega)-W_{J}(\omega^{\prime})\rVert_{J}^{2}\bigr{)}$ (9)
@@
For a fixed gridpoint $s_{m}\in S_{M}$ and given a state vector $\omega_{m}$ and the full training dataset $\{\omega_{nm}\}_{nm=0}^{NM-1}$, kernel $\kappa_{c}$ is used to produce a function $f_{m}\in H_{\Omega,N}$, $f_{m}(\cdot)=\kappa_{c}(\omega_{m},\cdot)$ leading to multiplication operator $M_{f_{m}}$ and the desired quantum effect $e_{m}=\operatorname{Proj}_{L}M_{f_{m}}$. Employing the quantum Bayes rule (2) then yields the conditioned quantum state $\rho(s_{m})|_{e_{m}}$ for gridpoint $s_{m}$. Intuitively, function $f_{m}$ can be thought of as an affinity function on state space $\Omega$ measuring how “similar” the given state $\omega_{m}$ is to the states included in our training dataset. Building operator $e_{m}$ and conditioning the prior state $\rho(s_{m})$ then yields the posterior state $\rho(s_{m})|_{e_{m}}$ that encodes the state space affinities captured by $f_{m}$. We often refer to the mapping $\omega_{m}\mapsto e_{m}$ as the feature map (as in Section 3.2) and to $f_{m}$ as the feature vector produced by kernel $\kappa_{c}$ from product state vector $\omega_{m}$.
With respect to our basis representation of $\rho(s_{m})$ by vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$ and effect $e_{m}$ by matrix $\bm{e}_{m}\in\mathbb{R}^{L\times L}$ the conditioning operation takes the form
@@
To avoid having to compute the square root of matrix $\bm{e}_{m}$ we modify the way we form effect $e_{m}$. We use function $f_{m}^{1/2}\in H_{\Omega,N}$ to form multiplication operator $\tilde{A}_{m}=M_{f_{m}}^{1/2}$ and define effect $\tilde{e}_{m}=(\operatorname{Proj}_{L}\tilde{A}_{m})^{2}=(\tilde{A}_{m}^{L})^{2}$. Conditioning with respect to $\tilde{e}_{m}$ now takes the form
@@
Although effects $\tilde{e}_{m}$ and $e_{m}$ are generally different, the two are equal when $L=N$ and thus also share the same asymptotic limit as $N\rightarrow\infty$ and $L\rightarrow\infty$ [55].
The data map $W_{J}$ embeds state vector $(x,s_{m})\in\Omega$ into $\mathbb{R}^{J}$ by considering the resolved state values at gridpoints neighboring $s_{m}\in S_{M}$, mimicking the structure of the stencil used for the spatial discretization of the considered dynamics. In this way kernel $k_{c}$ compares the stencil of points around $s_{m}$ to the corresponding stencil of points for every state included in our training data. This makes for a stronger affinity condition than comparing values of states at only a single gridpoint. For hyperbolic partial differential equations, where the notion of a local propagation velocity is well defined, using a stencil of values is analogous to using a time delayed signal.
Because the spatially local value $\omega_{m}=(x,s_{m})$ is compared to every other state vector present in our training data $\{\omega_{nm}\}_{nm=0}^{NM-1}$ spatial translations of state $x$ are naturally factored during the conditioning process. This means that if a spatial translation of state $x$ is present in our training dataset then the similarity between product state $\omega_{m}=(x,s_{m})$ and the corresponding value of the spatial translation of state $x$ will be identified in the conditioning process. For partial differential equations with spatial translation symmetry, this means that we do not have to artificially inflate our training dataset by including all spatial translations of the sampled states.
@@ 4.5 Multitrajectory data
To build the data driven formulation of our scheme we have so far relied on a single dynamical trajectory $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$. In this section we generalize this construction to the case where our training data consists of multiple trajectories. We begin by motivating the use of multiple trajectories before presenting the modifications required in the theoretical and data driven formulations of the scheme.
@@
Using a single dynamical trajectory is generally sufficient when considering a dynamical system possessing a unique physical measure, namely a measure that can be sampled by a set of initial conditions of positive measure with respect to the ambient measure [77]. These are measures whose sampling can be realized numerically, with SRB measures being an important example [78]. In such systems, a single long trajectory can be used to sample the physical measure of the dynamics.
@@
Using multiple trajectories as training data is of interest when considering dynamical systems possessing multiple physical measures; for instance, hamiltonian systems or hyperbolic systems with several competing basins of attraction. Additionally, using multiple trajectories is necessary in the forecasting of parameter dependent systems, where one considers a dynamical system whose qualitative dynamics changes with respect to a bifurcation parameter.
@@
Finally, even in systems with a unique physical measure, being able to use multiple trajectories to sample that measure allows one to consider optimal sampling strategies. In designing such a strategy one tries to optimize the sampling rate with respect to a chosen objective, instead of relying on the inherent rate provided by a single dynamical trajectory.
We consider a finite collection of invariant probability measures $\{\mu_{i}\}_{i}$ on $X$ with mutually disjoint supports where $i\in\{0,\ldots,I-1\}$ and $I\in\mathbb{N}$. We define the invariant probability measure $\mu=\sum_{i=0}^{I-1}\mu_{i}/I$ and form our quantum mechanical Hilbert space $H_{\Omega}=L^{2}(\Omega,\mu\times\nu_{M};\mathbb{C})$. Denoting $H_{\Omega,i}=L^{2}(\Omega,\mu_{i}\times\nu_{M};\mathbb{C})$ it follows that $H_{\Omega}$ admits the orthogonal decomposition
$H_{\Omega}\simeq\oplus_{i=0}^{I-1}H_{\Omega,i}$
[…]
To realize the above construction numerically we assume access to sampled trajectories $\{x_{n}^{(i)}\}_{n=0}^{N_{i}-1}$ of $N_{i}\in\mathbb{N}$ samples each, where $i\in\{0,\ldots,I-1\}$ as above. Similarly to the single trajectory case from Section 4.4, for each trajectory we define the empirical sampling measure $\mu_{i,N_{i}}=\sum_{n=0}^{N_{i}-1}\delta_{x_{n}^{(i)}}/N_{i}$ and assume that $\mu_{i,N_{i}}\to\mu_{i}$ in the weak-* sense as $N_{i}\to\infty$. We also define the empirical measure $\mu_{N}=\sum_{i=0}^{I-1}\mu_{i,N_{i}}/I$ and associated Hilbert space $H_{\Omega,N}=L^{2}(\Omega,\mu_{N}\times\nu_{M};\mathbb{C})$ with $N=\sum_{i=0}^{I-1}N_{i}$. By our assumption that the measures $\mu_{i}$ have mutually disjoint supports it follows that the sampled trajectories $\{x_{n}^{(i)}\}_{n=0}^{N_{i}-1}$ are also mutually disjoint. As a result we have the orthogonal decomposition
$H_{\Omega,N}\simeq\oplus_{i=0}^{I-1}H_{\Omega,N_{i}}$
with $H_{\Omega,N_{i}}=L^{2}(\Omega,\mu_{i,N_{i}}\times\nu_{M};\mathbb{C})$. Each component space $H_{\Omega,N_{i}}$ is a space of equivalence classes of functions represented by their values on the samples of the $i$th trajectory only. Leveraging the decomposition of $H_{\Omega,N}$ we use the approach described in Section 4.4 to build an eigenfunction basis for each component space $H_{\Omega,N_{i}}$ separately and use the $L_{i}\in\mathbb{N}$ leading eigenfunctions $\{\phi_{\ell}^{(i)}\}_{\ell=0}^{L_{i}-1}$ to form the subspace $H_{\Omega,N_{i}}^{L_{i}}\subset H_{\Omega,N_{i}}$. Denoting $L=\sum_{i=0}^{I-1}L_{i}$ we define the combined space
$H_{\Omega,N}^{L}=H_{\Omega,N_{0}}^{L_{0}}\oplus\cdots\oplus H_{\Omega,N_{I-1}}^{L_{I-1}}$
with the direct sum of the component bases producing a basis for the combined space $H_{\Omega,N}^{L}$.
The basis of $H_{\Omega,N}^{L}$ derived in this way consists of a total of $L$ elements, each being an eigenfunction of a block diagonal operator $K_{N}=K_{N_{0}}\oplus\ldots\oplus K_{N_{I-1}}$ with kernel integral operators $K_{N_{i}}\colon H_{\Omega,N_{i}}\to H_{\Omega,N_{i}}$ in its diagonal blocks. More specifically, each basis vector $\phi_{\ell}^{(i)}$ represents the values of a function on $N$ samples, with all of its values being zero except for the block of values that corresponds to the $N_{i}$ samples of the $i$th trajectory it was derived from. This approach has lower computational cost compared to building an eigenfunction basis for $H_{\Omega,N}$ by using a single kernel integral operator, which would require solving an eigenvalue problem for an $N\times N$ matrix. However, our approach also requires that each invariant measure $\mu_{i}$ is individually adequately sampled, as opposed to adequate joint sampling of measure $\mu$. This can potentially limit the number of high quality basis functions that can be derived from a dataset of fixed size $N$, for instance if one chooses to use a large number of trajectories with a small number of samples $N_{i}$ for each one.
Having derived the basis for space $H_{\Omega,N}^{L}$ the construction of the rest of the scheme proceeds as in Section 4.4. The only additional difference is in the definition of the shift operator $P_{N}$ where time shifts are now applied only within each trajectory, and the first temporal sample of each trajectory is mapped to zero. Namely, we make sure not to shift samples across different trajectories.
After forming our QMCl closure model in $H_{\Omega,N}^{L}$ our goal is to use it to predict the dynamics of a chosen dynamical system for out of sample initial conditions; namely, initial conditions lying outside of the basins of the measures $\mu_{i}$ sampled in the training stage. Thanks to the fact that the feature maps used to form quantum effects (Section 4.4) are not restricted to our training dataset, our closure scheme can ingest information from the resolved variables and render estimates of the fluxes even for out of sample data. Whether these estimates are sufficiently accurate ultimately depends on whether the kernel used for the conditioning of quantum states can identify samples of our training dataset whose corresponding flux evolution is similar to that of out of sample initial conditions. In the next section we present numerical results offering empirical evidence that out of sample predictions can indeed be made using QMCl for the tested dynamical system and family of initial conditions.